# A Unified, Practical, and Understandable Summary of Non-transactional Consistency Levels in Distributed Replication

### v1

@@ 4.5.7. Weak Consistency
Weak consistency is at the bottom of the consistency level spectrum and is weaker than all other consistency levels. In our model, weak consistency can be defined as enforcing an NPO and None-relationship ordering. It can simply be interpreted as “providing no consistency guarantees at all”. Note that this level is irrelevant to weak ordering in shared memory consistency (Mosberger, 1993; Hill, 1998).
@@ 4.5.8. Mixed/Hierarchical Consistency Levels
So far, we have assumed a single conceptual storage service without making any assumptions on the internal implementation of the service. Real distributed systems may, however, contain multiple layers or scopes of sub-services, each providing a different consistency level semantic. For example, CosmosDB (Microsoft, 2022) provides a stronger consistency guarantee for clients within the same region than those distributed across multiple regions, effectively exposing a 2-layer consistency model. Given the implementation details of a system, we can always refine the consistency level spectrum and define mixed or hierarchical consistency levels that are composed of multiple basic levels.
@@ 4.5.9. Memory Consistency Models
Distributed replication consistency is tightly related to early works in multiprocessor shared memory consistency. Hill defined hardware memory consistency model as the interface contract for memory in a shared memory multiprocessor, where instructions may be executed out-of-order (Hill, 1998). Many of the consistency levels described in this paper originated from multiprocessor shared memory consistency models. The three primary examples are sequential consistency, which was first defined by Lamport in (Lamport, 1979), causal consistency, which was adapted from causal memory (Ahamad et al., 1995), and PRAM consistency, which was adapted from FIFO processor consistency (Ahamad et al., 1993). Other relaxed memory consistency models and techniques, such as weak ordering, acquire/release consistency, entry consistency, cache coherence, and memory fences/barriers (Mosberger, 1993; Hill, 1998), are out of the scope of this paper.
@@ 5. Availability Guarantees
Besides consistency, availability is also an important part of the interface contract between a distributed storage service and clients. Availability is not implementation-oblivious; the meaning of fault-tolerance and availability can only be defined given a specific system model. In this section, we consider a simple system model of symmetric replicas and analyze the best possible availability guarantee that each consistency level can provide in such a system.
@@ 5.1. Symmetric Replicas System Model
We consider a fault-tolerant system implementation of the object store service composed of a set of symmetric replica servers, similar to what Figure 4 depicts. Each replica node holds a complete copy of all objects and can communicate with any other replica through message-passing over the network. Clients establish connections to one (or more) replica nodes, issue operations, and wait for acknowledgments.
Data Partitioning. Since we only consider non-transactional workloads, this symmetric model can be easily extended to incorporate data partitioning (or called partial replication), where each node is responsible for a subset of objects. For each object, only the set of nodes that hold the object is under consideration for availability of that object.
Client-side Caching. A client may act as a partial replica server by doing client-side coherent caching w.r.t. the consistency level for its reads and writes (Bailis et al., 2013c; Terry et al., 1994). In this case, we count the client itself as a valid partial replica. We leave a more detailed discussion on client-side caching as future work.
@@ 5.2. Meaning of Availability
We say a system of symmetric replicas provides high availability if, in the presence of arbitrarily long network partitions between arbitrary replicas, every client that can establish connection to one (or a specific set of) non-failing replica(s) of an object eventually gets acknowledged for all operations it issues on that object (Bailis et al., 2013a).
Failure Model. We allow the following two types of failures:
  • Fail-stop: a replica server may crash and stop responding to any incoming messages at any time point.
  • Network partition: the network channel between any two groups of replicas may break. A failed replica is indistinguishable to its peers from a completely network-partitioned replica or a replica that is just running very slowly.
We do not consider Byzantine faults (Castro et al., 1999; Mahajan et al., 2012; Fischer et al., 1985), where a server node may maliciously send out inconsistent information, or where the content of network messages may get tampered with. We leave the integration of Byzantine faults as future work.
Availability Levels. We consider three coarsely-defined levels of availability guarantees in the presence of arbitrarily long network partitions between arbitrary pairs of replicas (JEPSEN, 2016):
@@
  • Sticky available: a client maintains stickiness if it keeps contacting the same replica for each of its operations on an object. The system is sticky available if every client that sticks to a non-failing replica of an object eventually receives responses that honor the consistency level for operations on that object.
@@
Note that the “weakly available” category can be further decomposed into finer-grained, protocol-specific availability levels if we can bound the number of failures to a certain quantity. For example, most state machine replication protocols are available when at least a majority of nodes are healthy and can communicate with each other. Also, extra care needs to be taken to define reasonable transactional availability guarantees (Bailis et al., 2013a), which is out of the scope of this paper.
@@ 5.3. Availability Upper Bounds
Brewer’s famous CAP theorem states that a distributed system cannot achieve Consistency, Availability, and network Partition-tolerance all at the same time (Brewer, 2000). This informal description is often misunderstood and taken in an overly restrictive form. A more precise statement would be that a distributed system cannot achieve linearizability, total/sticky availability, and network partition-tolerance all at the same time. This statement has been proven by Gilbert and Lynch in (Gilbert and Lynch, 2002).
@@
Limitations. The availability upper bounds presented here are rather coarse-grained and do not capture everything about availability. First, they say nothing about recency guarantees, i.e., how stale are read results allowed to be. For example, although causal consistency is sticky available, a network partition may indefinitely prevent writes made on one side from being visible to readers on the other side. Bounded staleness levels would thus all be weakly available in our definition. Second, these availability bounds also do not consider partial network partitions, where certain pairs nodes cannot directly communicate with each other, but some indirect multi-hop paths are still available. Alfatafta et al. discussed partial network partitions and mechanisms to exploit indirect paths (Alfatafta et al., 2020).
@@ 6. Conclusion
This paper presents a unified, practical, and understandable summary of non-transactional consistency levels in the context of distributed data replication systems. We develop the intuitive shared object pool (SOP) model and define consistency levels within this framework by mapping them to two types of ordering validity constraints. We explain the four most common levels, namely linearizability, sequential, causal+, and eventual consistency, along with other refined levels with detailed examples. We also discuss the availability upper bound of presented levels.
As replicated, fault-tolerant systems become the norm in the cloud era, we hope this paper provides useful guidance for replication protocol designers and distributed system engineers.

### v2

@@ 4.5.7. Weak Consistency
Weak consistency [^7] Weak consistency is irrelevant to weak ordering in shared memory systems (Mosberger, 1993; Hill, 1998). is at the bottom of the consistency level spectrum and is weaker than all other consistency levels. In our model, weak consistency can be defined as enforcing an NPO and None-relationship ordering. It can simply be interpreted as “providing no consistency guarantees at all”.
@@ 4.5.8. Mixed/Hierarchical Consistency Levels
So far, we have assumed a single conceptual storage service without making any assumptions on the internal implementation of the service. Real distributed systems may, however, contain multiple layers or scopes of sub-services, each providing a different consistency level semantic. For example, CosmosDB (Microsoft, 2022) provides a stronger consistency guarantee for clients within the same region than those distributed across multiple regions, effectively exposing a 2-layer consistency model. Given the implementation details of a system, we can always define mixed or hierarchical consistency levels composed of multiple basic levels.
@@ 4.5.9. Memory Consistency Models
Distributed replication consistency is tightly related to early works in multiprocessor shared memory consistency. Hill defined hardware memory consistency model as the interface contract for shared memory, where instructions may be executed out-of-order (Hill, 1998). Memory consistency models and techniques such as weak ordering, acquire/release consistency, entry consistency, cache coherence, and memory fences/barriers (Mosberger, 1993; Hill, 1998) are out of the scope of this paper.
@@ 5. Availability Guarantees
Besides consistency, availability is also an important part of the interface contract between a distributed storage service and clients. Availability is not implementation-oblivious; the meaning of fault-tolerance and availability can only be defined given a specific system model. In this section, we consider a simple system of symmetric replicas and analyze the best possible availability guarantee that each consistency level can provide in such a system.
@@ 5.1. Symmetric Replicas System Model
We consider a fault-tolerant system implementation of the object store service composed of a set of symmetric replica servers, similar to what Figure 4 depicts. Each replica node holds a complete copy of all objects and can communicate with any other replica through messages over the network. Clients establish connections to one (or more) replica(s), issue operations, and wait for acknowledgments.
Data Partitioning. Since we only consider non-transactional workloads, this symmetric model can be easily extended to incorporate data partitioning (or called partial replication), where each node is responsible for a subset of objects. For each object, only the set of nodes that hold the object is under consideration for availability.
Client-side Caching. A client may act as a partial replica server by doing client-side coherent caching w.r.t. the consistency level for its reads and writes (Bailis et al., 2013b; Terry et al., 1994). In this case, we count the client itself as a valid partial replica.
@@ 5.2. Meaning of Availability
Consider a non-Byzantine fail-stop setting with an asynchronous network (Lamport, 2001). We say a system of symmetric replicas provides availability if, in the presence of arbitrarily long network partitions between arbitrary replicas, every client that can connect to one (or a specific set of) non-failing replica(s) of an object can get valid acknowledgments for all operations it issues on that object.
Availability Levels. We consider three coarsely-defined levels (JEPSEN, 2016):
@@
  • Sticky available: a client maintains stickiness if it keeps contacting the same replica for all of its operations on an object. The system is sticky available if every client that sticks to a non-failing replica of an object eventually receives responses that honor the consistency level for operations on that object.
@@
Note that the “weakly available” category can be further decomposed into finer-grained, protocol-specific availability levels if we can bound the number of failures to a certain quantity. For example, most state machine replication protocols are available when at least a majority of nodes are healthy and connected. Also, extra care needs to be taken to define reasonable transactional availability guarantees (Bailis et al., 2013a), which is out of the scope of this paper.
@@ 5.3. Availability Upper Bounds
The CAP theorem states that a distributed system cannot achieve Consistency, Availability, and network Partition-tolerance all at the same time (Brewer, 2000). This informal description is often taken in an overly restrictive form. A more precise statement would be that a distributed system cannot achieve linearizability, total/sticky availability, and tolerance to full network partitioning all at the same time. This statement has been proven by Gilbert and Lynch (Gilbert and Lynch, 2002).
@@
Limitations. The availability upper bounds presented here are rather coarse-grained and do not capture everything about availability. First, they say nothing about recency guarantees, i.e., how stale are read results allowed to be. For example, although causal consistency is sticky available, a network partition may indefinitely prevent writes made on one side from being visible to readers on the other side. Bounded staleness levels would thus all be weakly available in our definition. Second, these availability bounds do not consider partial network partitions, where certain pairs of nodes cannot directly communicate with each other, but some indirect multi-hop paths are still available. Alfatafta et al. discussed partial network partitions and mechanisms to exploit indirect paths (Alfatafta et al., 2020).
@@ 6. Conclusion
This paper presents a unified, practical, and understandable summary of non-transactional consistency levels in the context of distributed data replication systems. We develop an intuitive shared object pool (SOP) model and define consistency levels within this framework by constructing them out of two types of ordering validity constraints: convergence and relationship. We explain the four most common levels, namely linearizability, sequential, causal+, and eventual consistency, along with other refined levels with detailed examples. We also discuss their availability upper bound.
As replicated, fault-tolerant object storage systems become the cloud-era norm, we believe this paper provides useful guidance for replication protocol designers and distributed system engineers.